OpenStudy (anonymous):

Jacob bought eleven bottles of water. Each bottle contains one-half liter of water. How many liters did he buy? two and one-tenth five and one-fifth five and one-half eleven and one-half ill give medals help plz only the smartest people

OpenStudy (anonymous):

Knowing that (1/2+1/2 ) = 1 You could just add them. Which would give you: (1/2+1/2)+(1/2+1/2)+(1/2+1/2)+(1/2+1/2)+(1/2+1/2)+1/2 = 5 1/2 Know that adding the same number over and over again is just multiplication. Which would give you 11 X 1/2 = 11/1 X 1/2 = 11/2 Change the improper fraction to a mixed number by dividing the bottom number into the top number. So you have \[11\div2 = 5\frac{ 1 }{ 2}\]